OpenStudy (anonymous):

HELP MEDALS A square pyramid has a slant height of 13 meters and a height of 12 meters. What is the surface area if the square base has sides of length 10 meters? Answer 120 m2 260 m2 320 m2 360 m2

OpenStudy (anonymous):



OpenStudy (anonymous):

These are the triangle sides of the pyramid. The area is 13*10/2=65 There are 4 of these: 4*65=260 and the square base 10*10=100 So the answer is 360 m^2